OpenStudy (anonymous):

-3 + m / 9 = 10 A. -30 B. 63 C. 87 D. 93

HanAkoSolo (jamierox4ev3r):

-3+\(\frac{m}{9}=10\)

HanAkoSolo (jamierox4ev3r):

is this correct? @renae_ ?

OpenStudy (anonymous):

\[-3+\frac{ m }{ 9 }=10\]\[-3+3+\frac{ m }{ 9 }=10+3\]\[\frac{ m }{ 9 }=13\]\[m=9*13\]

OpenStudy (anonymous):

;)

HanAkoSolo (jamierox4ev3r):

-3+(9)\(\frac{m}{9}=10(9)\) -3+m=90 +3 +3 ---------- m=93

OpenStudy (anonymous):

no its, -3+m divided by 9 = 10 . all of -3+m is divided by 9

HanAkoSolo (jamierox4ev3r):

hmmm i got 93 @itsmylife ?

OpenStudy (anonymous):

incorrect babe ;)

OpenStudy (anonymous):

117

HanAkoSolo (jamierox4ev3r):

ohhhhh haha never mind :P i misunderstood the question.... Babe? lol since when

HanAkoSolo (jamierox4ev3r):

but thats not a choice >.<

OpenStudy (aonz):

\[ \frac{ -3+m }{ 9 } =10 \]

OpenStudy (anonymous):

m/9 = 10+3 m=9*13 m=117

OpenStudy (anonymous):

what @AonZ said

HanAkoSolo (jamierox4ev3r):

\(\frac{-3+m}{9}=10\)

OpenStudy (aonz):

yea therefore it is 93 liek @Jamierox4ev3r said

OpenStudy (anonymous):

117 is the right answer @Jamierox4ev3r

OpenStudy (johnweldon1993):

Right...93 as @Jamierox4ev3r has shown...

OpenStudy (anonymous):

i think you might have made mistake while writing your question

HanAkoSolo (jamierox4ev3r):

yay so i am correct i think.... idk now lol but 117 is not a choice and refrain from calling me babe plz lol :P

OpenStudy (anonymous):

but if your question is \[\frac{ -3+m }{ 9 }=10\] then 93 is correct

HanAkoSolo (jamierox4ev3r):

thank @johnweldon1993 i was beginning to question my abilities >.<

OpenStudy (anonymous):

yup, refrained ;)

OpenStudy (johnweldon1993):

haha nope...they're still sharp! xD